Roy services (Oil Change) vehicles in his garage. His lubricants cost him $40 per vehicle. His wages are $120 per hour. He spends 20 minutes per vehicle. His garage rent is $600 per month. He charges $140 to service each vehicle. His Contribution Ratio in servicing each vehicle is
Answer #1

Selling price per unit = $140

Lubricant cost per vehicle = $40

Wage rate = $120 per hour

Time spent per vehicle = 20 minutes

= 20/60 hours

= 1/3 hours

Wages per vehicle = Time spent per vehicle x Wage rate

= 120 x 1/3

= $40

Variable cost per unit = Lubricant cost per vehicle + Wages per vehicle

= 40 + 40

= $80

Fixed cost = Garage rent = $600

Contribution margin per unit = Selling price per unit – Variable cost per unit

=140-80

= $60

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 60/140

= 42.8571429%

Contribution Ratio in servicing each vehicle is 42.8571429%.

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!